need help with solving B please and give explanation
media%2F134%2F13415c88-4336-4196-96e3-62
Answer #1

(b) Given that, mass of an object = 14.01 g

Volume of water alone in beaker = 25.2 mL

Volume of water and object in beaker = 33.6 mL

Volume of object = Volume of water and object in beaker - Volume of water alone in beaker = 33.6 - 25.2 = 8.4 mL

Now, Density = mass / Volume = 14.01 g / 8.4 mL = 1.67 g/mL